OpenStudy (lynn):

Algegra: If 4/5 X ____=2/5 what is _____

OpenStudy (anonymous):

4 2 - x = -- 5 5 multiply both sides by the reciprocal 5/4 5 4 2 5 - * - x = - * - 4 5 5 4 x = 10 -- 20 x = 1/2

OpenStudy (lynn):

Thank you my answers were all wrong